OpenStudy (anonymous):

simplify z=(-2xy^2-y^3)/(3xy^2+2x^2y)

OpenStudy (anonymous):

the answer is (-2xy-y^2)/(3xy+2x^2) but how come you can't simplify this no more if there's an xy on top and bottom?

hartnn (hartnn):

because u cannot cancel xy like that \[\frac{a+b}{a+c}\neq \frac{b}{c}\]

OpenStudy (anonymous):

then how come you were able to cancel out from the orig

hartnn (hartnn):

from original,u only took y common from numerator and denominator,then cancelled,its like this: \[\frac{ab+ac}{ad+ae}=\frac{a(b+c)}{a(d+e)}=\frac{b+c}{d+e}\]

OpenStudy (anonymous):

omg thank you again summer made me forget math:P

hartnn (hartnn):

welcome :)